on a hypothetical planet X with an unknown value for g, the period of a swinging pendulum with a string length of .75 meters is a 1.8 s calculate the value of "g" on plnaet X



1
Expert's answer
2021-12-17T11:50:50-0500

The period of the pendulum is given as follows:


"T = 2\pi \sqrt{\dfrac{L}{g}}"

where "L = 0.75m" is its length. Thus, expressing "g" and substituting "T = 1.8s", find:


"g = \dfrac{4\pi^2L}{T^2} =\dfrac{4\pi^2\cdot 0.75m}{(1.8s)^2} \approx 9.1m/s^2"

Answer. 9.1 m/s^2.